Fixed cost is 400, firm's average total cost is 3 , and average variable cost is 2.50 it's output?
1
Expert's answer
2021-08-17T16:56:47-0400

"AFC=\frac{FC}{Q}"


"=\frac{400}{Q}"


"ATC=AFC+AVC"

"3=\frac{400}{Q}+2.50"

"Q=\frac{400}{0.50}"


"Q=800"